I'm having problems putting this into an equation...
One number is five greater than another. If the sum of the lesser number and twice the greater is 61, what is the greater number?
My Work:
I don't even know where to start....
Thanks!

Found 2 solutions by venugopalramana, pwac:
Answer by venugopalramana(3286)

One number is five greater than another. If the sum of the lesser number and twice the greater is 61, what is the greater number?
let the smaller number be x
bigger number is more by 5 that is x+5
twice the greater number =2*(x+5)=2x+10
sum it with lesser number ,we get =2x+10+x=3x+10
this sum is 61..so 3x+10=61
3x=61-10=51
x=51/3=17
so smaller number =17
bigger number is x+5=17+5=22

Answer by pwac(253)

lesser number = x
greater number = x+5 so twice greater number=2(x+5)
equation is:
x+2(x+5)=61
x+2x+10=61
3x=51
x=17 lesser number
greater number is 17+5=22 then doubled=44
17+44=61
Pete